OpenStudy (anonymous):

Let v_1, v_2 and v_3 be linearly independent vectors in a vector space V . Prove that w_1 = 2*v_1, w_2 = 2*v_1 + v_2 + v_3 and w_3 = v_1 + v_2 - v_3 are also linearly independent... Help? Please? :)

OpenStudy (zarkon):

suppose it is not L.I. and draw a contadiction

OpenStudy (zarkon):

or use the contrapositive

OpenStudy (anonymous):

contadicion? lol you seem to be the one to answer these questions of mine every time! i appreciate the patience :P

OpenStudy (zarkon):

assume \[c_1w_1+c_2w_2+c_3w_3=0 \] but at least on of the \[c_i\neq 0\]

OpenStudy (zarkon):

expand using the definitions of w1,w2,w3

OpenStudy (zarkon):

hopefully this will work...I'm just thinking out loud

OpenStudy (zarkon):

\[c_12v_1+c_2(2v_1 + v_2 + v_3)+c_2( v_1 + v_2 - v_3)=0 \]

OpenStudy (zarkon):

\[(2c_1+2c_2+c_3)v_1+(c_2+c_3)v_2+(c_2-c_3)v_2=0\]

OpenStudy (zarkon):

we have to have \[(2c_1+2c_2+c_3)=(c_2+c_3)=(c_2-c_3)=0\] since the v's are LI

OpenStudy (zarkon):

if we solve this system for the c's we get all the c's are zero...a contradiction

OpenStudy (zarkon):

ok?

OpenStudy (anonymous):

Yep - I get all of the maths (thanks!) ... But I might be missing the essential definition of LI .. why is it a contradiction??

OpenStudy (zarkon):

we originally assumed at least one of the c's was non-zero...then we get that they are all zero

OpenStudy (anonymous):

Why do we make that initial assumption?:/

OpenStudy (zarkon):

because I'm trying to prove it by contradiction the def of LI says that the vectors \[v_1,v_2,\ldots,v_n\] are LI iff \[c_1v_1+c_2v_2+\ldots+c_nv_n=0\] has only the trivial solution

jimthompson5910 (jim_thompson5910):

He's saying that the set of vectors {v1,...,vn} is linearly independent iff all the coefficients are 0 So assume that the set is NOT independent, then prove that a contradiction arises because of this assumption. So the contradiction would then point you in the opposite direction.

OpenStudy (anonymous):

Onto it! Thanks guys!!

jimthompson5910 (jim_thompson5910):

The opposite direction being that they are in fact linearly independent